Solve a two-period storage problem in which production is 250 in the first period and 150 in the second, while demand in each period is 200-P1 and 220-P2, respectively. Assume storage cost is 5+0.1P1 for each unit stored. Make sure you find 1) both prices and 2) quantities sold in each period, and 3) the quantity stored.
Answer #1

Total Production - Productin . Productim Period 2 Total Production - 2 t 150 0 400 Total Production - Total Consumption - Con

Salving equation ① & ② P = 20-P2 - 0) P = 1.1 Pits (2) P = 1.1 (20-2) + S = 22 - 1.1 P₂ +5 2.1P = 27 - P2 = 12.85 Approx) = P

The quantity Stooed il calcolated as follows Storage en Period 1 = Production - Consumption = 250 - 192.85 I 97.15 1 tve) Sto